At the high school band's annual spaghetti dinner, the members served 210 dinners and collected $935. The band leader did not designate which ticket was an adult or child dinner. You need to know that to prepare the correct amount of food and for future planning. Each adult dinner cost $6 and each child's dinner cost $3.50.
I understand the equations
6a+3x5c = 935
a+c= 210
How do I figure out how many dinners were adults and how many were children's?

Answer by checkley71(8403)

A+C=210 OR A=210-C FOR THE TICKETS SOLD.
6A+3.5C=935 NOW REPLACE A WITH(210-C) IN THIS EQUATION & SOLVE FOR C
6(210-C)+3.5C=935
1260-6C+3.5C=935
-2.5C=935-1260
-2.5C=-325
C=-325/-2.5
C=130 CHILDREN TICKETS SOLD.
A=210-130
A=80 ADULT TICKETS SOLD.
PROOF
6*80+3.5*130=935
480+455=935
935=935